# Contextual Gesture: Co-Speech Gesture Video Generation through Context-aware Gesture Representation

### v1

@@ Appendix E Gesture Motion Generation
Ablation.
We evaluate the mask ratio during training and the number of inference steps during decoding. As shown in Tab. 3(b), our model requires only 5 inference steps, in contrast to over 50 or 100 steps in diffusion-based models. Furthermore, a uniform masking ratio between 0.5 and 1 during training yields optimal performance.
@@ Appendix F Speech-Gesture Retrieval
Evaluation Result. Shown in Tab. 4, the gesture patterns and speech context do not present one-to-one mapping relationship, leading to the significantly low performance of retrieval. However, based on setting (b), within a small batch size of 32, the model achieves significantly higher performance, indicating the alignment provides the discrimination over different speech context and the motion. In addition, chronological negative examples during contrastive training enhances the robustness of retrieval.
@@ Appendix G Additional Experiments
In the main paper, we have shown our method achieves promising joint gesture motion and video generation. To understand the disentangled gesture and video avatar generation separately, we further conduct Gesture Generation and Video Avatar Animation experiments separately to compare our method with the corresponding representative works for each domain.
| Methods | Venue | FGD | BC | Diversity | MSE |
| GT | 0.703 | 11.97 | |||
| HA2G (Liu et al., 2022d) | CVPR 2022 | 12.320 | 0.677 | 8.626 | - |
| DisCo (Liu et al., 2022a) | ACMMM 2022 | 9.417 | 0.643 | 9.912 | - |
| CaMN (Liu et al., 2022b) | ECCV 2022 | 6.644 | 0.676 | 10.86 | - |
| DiffSHEG (Chen et al., 2024) | CVPR 2024 | 7.141 | 0.743 | 8.21 | 9.571 |
| TalkShow (Yi et al., 2023) | CVPR 2023 | 6.209 | 0.695 | 13.47 | 7.791 |
| Rhythmic Gesticulator (Ao et al., 2022) | SIGGRAPH 2023 | 6.453 | 0.665 | 9.132 | |
| EMAGE (Liu et al., 2023) | CVPR 2024 | 5.512 | 0.772 | 13.06 | 7.680 |
| Ours (w/o distill) | - | 5.079 | 0.737 | 13.24 | 7.742 |
| Ours | - | 4.434 | 0.724 | 13.76 | 7.021 |
G.1 Gesture Generation
Experiment Settings We select BEAT-X (Liu et al., 2023) as the dataset for additional gesture generation comparison. For consistency, we will exclude the image-to-animation component from our method and extend gesture representation from 2D to 3D poses. (with SMPL-X expressions for face gestures, as in the existing literature) We compare the gesture generation module of our work with representative state-of-the-art methods in co-speech gesture generation (Ao et al., 2022; Yi et al., 2023; Liu et al., 2023). We further design a baseline without using contextual distillation.
Evaluation Metrics We evaluate the realism of body gestures by Fréchet Gesture Distance (FGD)(Yoon et al., 2020). We include Diversity by calculating the average L1 distance across generated clips. For synchronization, we use Beat Constancy (BC) (Li et al., 2021). For facial expression, we use the vertex Mean Squared Error (MSE) (Xing et al., 2023) to for positional accuracy.
Experiment Results As shown in Tab. 5, our method significantly improve the SMPL-X based co-speech gesture generation with lower FGD and higher diversity. Specifically, Our methods have present smoother gesture motion patterns compared with existing works. It demonstrates the effectiveness of contextual distillation for the motion representation learning in our framework. We defer the video comparisons in the Appendix videos for reference.
| Dataset | Setting | FGD | Diversity | BAS |
| PATS | 10s | 1.303 | 13.260 | 0.996 |
| 10s | 2.356 | 11.956 | 0.994 | |
| BEAT-X | 10s | 4.747 | 13.14 | 7.323 |
| 10s | 4.650 | 13.55 | 7.370 | |
Long Sequence generation To understand the capability of our framework for long sequence generation, we conduct an ablation study for both PATS and BEAT-X dataset. For BEAT-X, we cut the testing audios into segments of 256 (about 8.53 seconds) for short sequence evaluation and use raw testing audios for long sequence evaluation in Tab. 5. Shown in Tab. 6, it is interesting for PATS dataset, long-sequence generation as an application in the main paper presents quality lower than normal settings. However, for BEAT-X dataset, the generation quality is not affected much. We attribute this difference caused by the dataset difference. Because PATS dataset consists training video lengths with a average of less than 10 seconds, the model presents less diverse gesture patterns. However, in BEAT-X, most of gesture video sequences are over 30 or 1 minutes, our method further benefits from this long sequence learning precess and presents higher qualities.
G.2 Video Avatar Animation
@@ Appendix H Time and Resource Consumption
In Tab. 7, we present a comparison of training and inference times against existing baseline methods. For audio-gesture generation, our model’s training time is comparable, albeit slightly slower, than that of ANGIE (Liu et al., 2022c) and S2G-Diffusion (He et al., 2024), primarily due to the inclusion of additional modules. However, it is considerably faster than MM-Diffusion (Ruan et al., 2023). Notably, our method excels in inference speed, outperforming all other baselines.
@@
We further compared image-warping based method computation requirements with Stable Diffusion-based models like AnimateAnyone (Hu et al., 2023) in Tab. 8.
@@
| Ours | 6d | Quantize 0.2d + Mask-Gen 1.5d + Res-Gen 0.5d + Img-warp & Refine 3.5d | 3 sec | ||
@@ Appendix I User Study Details
For user study, we recruited 20 participants with good English proficiency. To conduct the user study, we randomly select 80 videos from EchoMimicV2 (Meng et al., 2024), ANGIE (Liu et al., 2022c), S2G-Diffusion (He et al., 2024) and ours. Each user works on 20 videos, with 4 videos from each of the aforementioned methods. The users are not informed of the source of the video for fair evaluations. A visualization of the user study is shown in Fig. 9.
@@
In this paper, we utilize Thin Plate Splines (TPS) (Zhao & Zhang, 2022) to model deformations based on human poses for image-warping. Here, we provide additional details on this approach.
@@
where and denote the keypoints in and respectively. As shown in (Zhao & Zhang, 2022), it can be demonstrated that the TPS interpolating function satisfies Eq. 11:
where represents the coordinates in , and is the keypoint in . The function serves as a radial basis function. Notably, is the fundamental solution to the biharmonic equation (Selvadurai & Selvadurai, 2000), defined by:
@@
Although it is difficult to eliminate the potential for misuse entirely, our research offers a valuable technical analysis of gesture video synthesis. This contribution is intended to enhance understanding of the technology’s capabilities and limitations, particularly concerning details such as facial nuances and temporal coherence.

### v2

@@ Appendix F Speech-Gesture Retrieval
Evaluation Result. Shown in table 4, the gesture patterns and speech context do not present one-to-one mapping relationship, leading to the significantly low performance of retrieval. However, based on setting (b), within a small batch size of 32, the model achieves significantly higher performance, indicating the alignment provides the discrimination over different speech context and the motion. In addition, chronological negative examples during contrastive training enhances the robustness of retrieval.
@@ Appendix G Additional Experiments
In the main paper, we have shown our method achieves promising joint gesture motion and video generation. To understand the disentangled gesture and video avatar generation separately, we further conduct Video Avatar Animation experiments separately to compare our method with the corresponding representative works.
G.1. Video Avatar Animation
@@ Appendix H Time and Resource Consumption
In table 5, we present a comparison of training and inference times against existing baseline methods. For audio-gesture generation, our model's training time is comparable, albeit slightly slower, than that of ANGIE (Liu et al., 2022b) and S2G-Diffusion (He et al., 2024), primarily due to the inclusion of additional modules. However, it is considerably faster than MM-Diffusion (Ruan et al., 2023). Notably, our method excels in inference speed, outperforming all other baselines.
@@
We further compared image-warping based method computation requirements with Stable Diffusion-based models like AnimateAnyone (Hu et al., 2023) in table 6.
@@
| Ours | 6d | Quantize 0.2d + Mask-Gen 1.5d + Res-Gen 0.5d + Img-warp & Refine 3.5d | 3 sec |
@@ Appendix I User Study Details
For user study, we recruited 20 participants with good English proficiency. To conduct the user study, we randomly select 80 videos from EchoMimicV2 (Meng et al., 2024), ANGIE (Liu et al., 2022b), S2G-Diffusion (He et al., 2024) and ours. Each user works on 20 videos, with 4 videos from each of the aforementioned methods. The users are not informed of the source of the video for fair evaluations. A visualization of the user study is shown in fig. 9.
@@
In this paper, we utilize Thin Plate Splines (TPS) (Zhao and Zhang, 2022) to model deformations based on human poses for image-warping. Here, we provide additional details on this approach.
@@
where and denote the keypoints in and respectively. As shown in (Zhao and Zhang, 2022), it can be demonstrated that the TPS interpolating function satisfies eq. 11:
where represents the coordinates in , and is the keypoint in . The function serves as a radial basis function. Notably, is the fundamental solution to the biharmonic equation (Selvadurai and Selvadurai, 2000), defined by:
@@
Although it is difficult to eliminate the potential for misuse entirely, our research offers a valuable technical analysis of gesture video synthesis. This contribution is intended to enhance understanding of the technology's capabilities and limitations, particularly concerning details such as facial nuances and temporal coherence.